OpenStudy (anonymous):

If sin Θ = 12 over 37, use the Pythagorean Identity to find cos Θ. cos Θ = ± 35 over 37 cos Θ = ± 23 over 37 cos Θ = ± 35 over 12 cos Θ = ± 20 over 35

OpenStudy (anonymous):

y=12 and z=37\[x ^{2}+y ^{2}= z ^{2}\] Solve for x \[The \cos \theta = x/z\]